OpenStudy (anonymous):

Dan earns $42.75 per month working as a tutor. He worked as a tutor for 4 monthsHe also earned $52.45 walking dogs in the neighborhood. He spent $18.50 of his total earning to buy a movie ticket and candy. Dan put 1 over 3 of the remaining money in his savings account Dan used the following steps to find the total money he put in his savings account: Step 1: [($42.75 ⋅ 4) + $52.45 − $18.50] ⋅ 1 over 3 Step 2: [$171 + $52.45 − $18.50] ⋅ 1 over 3 Step 3: 241.95 ⋅ 1 over 3 Step 4: $80.65 In which step did Dan first make an error Use words to explain how Dan can correct his error

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

\(\large 42.75 \times 4=x\)

OpenStudy (anonymous):

171

OpenStudy (dangerousjesse):

\(\large 171+52.45=y\)

OpenStudy (anonymous):

223.45

OpenStudy (dangerousjesse):

\(\large 223.45-18.5=z\)

OpenStudy (dangerousjesse):

You should see which one he did wrong by now :)

OpenStudy (anonymous):

he went wrong at 223.45- 18.5

OpenStudy (dangerousjesse):

Yes :)